### v4

# BiSSL: Enhancing the Alignment Between Self-Supervised Pretraining and Downstream Fine-Tuning via Bilevel Optimization

Abstract
Models initialized from self-supervised pretraining may suffer from poor alignment with downstream tasks, reducing the extent to which subsequent fine-tuning can adapt pretrained features toward downstream objectives. To mitigate this, we introduce BiSSL, a novel bilevel training framework that enhances the alignment of self-supervised pretrained models with downstream tasks prior to fine-tuning. BiSSL acts as an intermediate training stage conducted after conventional self-supervised pretraining and is tasked with solving a bilevel optimization problem that incorporates the pretext and downstream training objectives in its lower- and upper-level objectives, respectively. This approach explicitly models the interdependence between the pretraining and fine-tuning stages within the conventional self-supervised learning pipeline, facilitating enhanced information sharing between them that ultimately leads to a model initialization better aligned with the downstream task. We propose a general training algorithm for BiSSL that is compatible with a broad range of pretext and downstream tasks. Using SimCLR and Bootstrap Your Own Latent to pretrain ResNet-50 backbones on the ImageNet dataset, we demonstrate that our proposed framework significantly improves accuracy on the vast majority of 12 downstream image classification datasets, as well as on object detection. Exploratory analyses alongside investigative experiments further provide compelling evidence that BiSSL enhances downstream alignment.
1 Introduction
In the absence of sufficient labeled data, self-supervised learning (SSL) has proven to be a powerful framework for training deep learning models. SSL leverages large amounts of unlabeled data to learn representations that can be effectively utilized across various downstream tasks. These representations are learned by solving a pretext task, which operates on supervisory signals extracted from the unlabeled data itself. Extensive efforts have been dedicated to designing effective pretext tasks, leading to state-of-the-art or competitive performance in various fields such as computer vision [1, 2, 3, 4, 5, 6, 7, 8], natural language processing [9, 10, 11, 12, 13] and audio signal processing [14, 15, 16, 17, 18, 19, 20]. A key advantage of SSL is that it enables the reuse of pretrained model backbones, eliminating the need to train models from scratch and thereby facilitating efficient adaptation to specific downstream tasks.
Figure 1: The conventional self-supervised learning pipeline alongside the proposed pipeline involving BiSSL. The symbols ${\bf\it\theta}$ and ${\bf\it\phi}$ represent backbone and task-specific attached head parameters, respectively. When they are transmitted to the respective subsequent training stages, they are used as initializations. The objectives $\mathcal{L}^{P}$, $\mathcal{L}^{D}$ represent the respective pretext pretraining and downstream fine-tuning objectives and $\mathcal{D}^{P}$, $\mathcal{D}^{D}$ the respective unlabeled pretext and labeled downstream datasets. We refer to Section 3 for further details.
When a model is pretrained on a dataset that distributionally differs from the downstream dataset, the representations learned during pretraining are likely to be misaligned with the downstream task. During fine-tuning [21, 22], this distribution misalignment is prone to causing relevant features, learned during the pretraining phase, to vanish from the representation space [23, 24, 25]. A strategy for mitigating the negative effects of these distribution discrepancies is to enhance the alignment between the self-supervised pretraining and downstream fine-tuning stages. However, since the conventional SSL pipeline treats these stages as two disjoint processes, this poses a significant challenge in devising a strategy that enhances such alignment while not compromising on the benefits that SSL offers.
We suggest utilizing bilevel optimization (BLO) as a novel approach for enhancing this alignment. BLO has risen as a powerful tool for solving certain optimization problems within deep learning. It entails a main optimization problem constrained by the solution to a secondary optimization problem that depends on the parameters of the main optimization objective. This hierarchical setup causes the solutions of both optimization problems to depend on each other, either directly or implicitly, which has proven advantageous in deep learning tasks that optimize multiple inter-dependent objectives simultaneously [26]. Notable mentions of tasks within deep learning where BLO has proven useful are parameter pruning [27], invariant risk minimization [28, 29], meta-learning [30, 31], adversarial robustness [32], hyper-parameter optimization [33] and coreset selection [34].
In this study, we propose BiSSL, a novel bilevel training framework that improves the alignment between the pretext pretraining and downstream fine-tuning stages in SSL. Acting as an intermediate training stage within the SSL pipeline, BiSSL frames the pretext and downstream task objectives as the lower- and upper-level objectives of a BLO problem. These objectives are connected by substituting the lower-level backbone solution into the upper-level objective. Simultaneously, the lower-level backbone is constrained to resemble the upper-level backbone parameters. This approach explicitly models the interdependence between pretext pretraining and downstream fine-tuning, enabling more effective information sharing between the pretext and downstream stages, yielding a backbone initialization better aligned with the downstream task. Figure 1 provides an overview that compares the conventional SSL pipeline with our proposed training pipeline involving BiSSL. We introduce a training algorithm that alternates between optimizing the two objectives within BiSSL, which is agnostic to the choice of pretext and downstream task. It is applied after the SSL pretraining stage, and imposes no additional assumptions about the pretraining procedure, allowing compatibility with off-the-shelf pretrained models. Through experiments using both the SimCLR [1] and Bootstrap Your Own Latent [4] pretext tasks to pretrain ResNet-50 backbones [35] on the ImageNet-1K dataset [36], we demonstrate that BiSSL significantly improves downstream performance across most of 12 downstream image classification datasets, while also significantly improving object detection accuracy. The code implementation and pretrained model weights are publicly available. [^1] https://github.com/NIPS25-16372/NIPS25_16372_BiSSL
2 Related Work
Bilevel Optimization in Self-Supervised Learning
Bilevel optimization (BLO) refers to a constrained optimization problem, where the constraint itself is a solution to another optimization problem that depends on the parameters of the “main” optimization problem. The general BLO problem is formulated as
$\underset{{\bf\it\xi}}{\min}\,f\quantity({\bf\it\xi},{\bf\it\psi}^{\ast}({\bf\it\xi}))\quad\text{s.t.}\quad{\bf\it\psi}^{\ast}({\bf\it\xi})\in\underset{{\bf\it\psi}}{\text{argmin}}\,g\quantity({\bf\it\xi},{\bf\it\psi}),$ (1)
where $f$ and $g$ are referred to as the upper-level and lower-level objectives, respectively. While the lower objective $g$ has knowledge of the parameters ${\bf\it\xi}$ from the upper-level objective, the upper-level objective $f$ possesses full information of the lower objective $g$ itself through its dependence on the lower-level solution ${\bf\it\psi}^{\ast}({\bf\it\xi})$. Some works have incorporated bilevel optimization within self-supervised learning. [37] suggest formulating the contrastive self-supervised pretext task as a bilevel optimization problem, dedicating the upper-level and lower-level objectives for updating the backbone and projection head parameters respectively. Other frameworks such as the Local and Global (LoGo) [38] and Only Self-Supervised Learning (OSSL) [39] utilize auxiliary models, wherein the lower-level objective optimizes the parameters of the auxiliary model, while the upper-level objective is dedicated to training the feature extraction model. MetaMask [40] introduces a meta-learning based approach, where the upper-level learns masks that filter out irrelevant information from inputs that are provided to a lower-level self-supervised contrastive pretext task. BLO-SAM [41] is tailored towards fine-tuning the segment anything model (SAM) [42] by interchangeably alternating between learning (upper-level) prompt embeddings and fine-tuning the (lower-level) segmentation model. The aforementioned frameworks integrate bilevel optimization into either the pretraining or fine-tuning stage exclusively and are tailored towards specific pretext or downstream tasks. In contrast, our proposed BiSSL employs a BLO problem that comprehensively incorporates both training stages of pretext pretraining and downstream fine-tuning, while not being confined to any specific type of pretext or downstream task.
Priming Pretrained Backbones Prior To Fine-Tuning
Previous works have demonstrated that downstream performance can be enhanced by introducing techniques that modify the backbone between the pretraining and fine-tuning stages. Contrastive Initialization (COIN) [43] introduces a supervised contrastive loss to be utilized on backbones pretrained with contrastive SSL techniques. Noisy-Tune [44] perturbs the pretrained backbone with tailored noise before fine-tuning. Speaker-invariant clustering (Spin) [45] utilizes speaker disentanglement and vector quantization for improving speech representations for speech signal specific downstream tasks. RIFLE [46] conducts multiple fine-tuning sessions sequentially, where the attached downstream specific layers are re-initialized in between every session. Unlike BiSSL, these techniques either do not incorporate knowledge of both the pretext task and downstream task objectives and their relationship, or do so only implicitly.
3 Proposed Method
3.1 Notation
We denote the unlabeled pretext dataset $\mathcal{D}^{P}=\{\mathbf{z}_{k}\}_{k=1}^{C_{P}}$ and labeled downstream dataset $\mathcal{D}^{D}=\{\mathbf{x}_{l},\mathbf{y}_{l}\}_{k=1}^{C_{D}}$, respectively, where $\mathbf{z}_{k},\mathbf{x}_{l}\in\mathbb{R}^{N}$. Let $f_{{\bf\it\theta}}:\mathbb{R}^{N}\to\mathbb{R}^{M}$ denote a feature extracting backbone with trainable parameters ${\bf\it\theta}$ and $h^{T}_{{\bf\it\phi}}:\mathbb{R}^{M}\to\mathbb{R}^{Q_{T}}$ a task specific head with trainable parameters ${\bf\it{\phi}}$. Given pretext and downstream models $h^{P}_{{\bf\it\phi}_{P}}\circ f_{{\bf\it\theta}_{P}}$ and $h^{D}_{{\bf\it\phi}_{D}}\circ f_{{\bf\it\theta}_{D}}$ with ${\bf\it\theta}_{P},{\bf\it\theta}_{D}\in\mathbb{R}^{L}$, we denote the pretext and downstream training objectives $\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi}_{P};\mathcal{D}^{P})$ and $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D};\mathcal{D}^{D})$, respectively. To simplify notation, we omit the dataset specification from the training objectives, e.g. $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D}):=\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D};\mathcal{D}^{D})$.
3.2 Optimization Problem Formulation
The conventional setup of self-supervised pretraining directly followed by supervised fine-tuning relies on using a single backbone model with parameters ${\bf\it\theta}$. In that instance, we minimize $\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi}_{P})$ to produce a backbone parameter configuration ${\bf\it\theta}_{P}^{*}$ which is then used as an initialization when subsequently minimizing the downstream training objective $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D})$. We deviate from this by instead considering ${\bf\it\theta}_{P}$ and ${\bf\it\theta}_{D}$ as two separate parameter vectors that are strongly correlated. In the aim of achieving an initialization ${\bf\it\theta}^{*}_{P}$ that is better aligned with the downstream task, we suggest an intermediate training stage that combines the two traditionally separate optimization problems of pretext and downstream training into a joint optimization problem through bilevel optimization, which we term BiSSL. The bilevel optimization problem is formulated as
$\underset{\boldsymbol{\theta}_{D},\boldsymbol{\phi}_{D}}{\min}\,\,\mathcal{L}^{D}\left(\boldsymbol{\theta}_{P}^{\ast}\left(\boldsymbol{\theta}_{D}\right),\boldsymbol{\phi}_{D}\right)+\mathcal{L}^{D}\left(\boldsymbol{\theta}_{D},\boldsymbol{\phi}_{D}\right)$ (2)
[…]
with $r$ being some convex regularization objective weighted by $\lambda\in\mathbb{R}_{+}$ enforcing similarity between ${\bf\it\theta}_{D}$ and ${\bf\it\theta}_{P}$. The upper-level training objective in (2) is tasked with minimizing the downstream task objective $\mathcal{L}^{D}$, while the lower-level objective in (3) aims to minimize the pretext task objective $\mathcal{L}^{P}$ while also ensuring its backbone remains similar to the upper-level backbone. As seen in the left term of (2), the backbone parameters ${\bf\it\theta}^{*}_{P}({\bf\it\theta}_{D})$ are transferred into the downstream training objective, mirroring how the backbone is transferred in the conventional SSL pipeline. The second term of (2) aids convergence of the upper-level optimization during training. It also plays a crucial role in ensuring that the lower-level optimization problem (3) is non-redundant and we refer the reader to Section A.1 of Appendix A for details. Unlike the traditional SSL setup, the backbone solution of the pretext objective ${\bf\it\theta}^{*}_{P}({\bf\it\theta}_{D})$ is now a function of the parameters of the downstream backbone ${\bf\it\theta}_{D}$, since the lower-level problem is dependent on the upper-level backbone parameters.
As the upper-level objective in (2) depends on the solution ${\bf\it\theta}^{*}_{P}({\bf\it\theta}_{D})$ of the lower-level problem in (3), this enables the incorporation of information from the pretext objective when solving the upper-level optimization problem. By including a regularization objective $r$ that enforces similarity between the lower-level and upper-level backbone parameters, this setup is expected to guide the lower-level optimization toward a configuration of model backbone parameters that is more beneficial for subsequent fine-tuning on the downstream task. To more specifically understand how the pretext objective influences the downstream training procedure in this setup, we delve deeper into the gradient expression of the upper-level training objective in (2) in the following subsection.
3.3 Upper-level Derivative
Given the upper-level objective $F({\bf\it\theta}_{D},{\bf\it\phi}_{D}):=\mathcal{L}^{D}({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\phi}_{D})+\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D})$ from (2), its derivative with respect to ${\bf\it\theta}_{D}$ is given by
$\begin{split}&\frac{\text{d}F}{\text{d}{\bf\it\theta}_{D}}=\underbrace{\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}}_{\text{IJ}}\nabla_{{\bf\it\theta}}\mathcal{L}^{D}\quantity({\bf\it\theta},{\bf\it\phi}_{D})|_{{\bf\it\theta}={\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}+\nabla_{{\bf\it\theta}}\mathcal{L}^{D}\quantity({\bf\it\theta},{\bf\it\phi}_{D})|_{{\bf\it\theta}={\bf\it\theta}_{D}}.\end{split}$ (4)
Due to the dependence of the lower-level solution on the upper-level parameters, the first term of (4) includes the jacobian of the implicit function ${\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})$, referred to as the implicit jacobian (IJ). To simplify notation, we let $\nabla_{{\bf\it\xi}}h({\bf\it\xi})|_{{\bf\it\xi}={\bf\it\psi}}:=\nabla_{{\bf\it\xi}}h({\bf\it\psi})$ when it is clear from context which variables are differentiated with respect to. Following an approach similar to [30], with details on the derivations and underlying assumptions outlined in Section A.2 of Appendix A, the IJ in (4) can be explicitly expressed as
$\begin{split}&\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}=-\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}r({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))\quantity[\nabla_{{\bf\it\theta}}^{2}\quantity(\frac{1}{\lambda}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+r\quantity({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})))]^{-1}.\end{split}$ (5)
A common convex regularization objective, which will also be the choice in the subsequent experiments of this work, is $r\quantity({\bf\it\xi},{\bf\it\psi})=\frac{1}{2}\norm{{\bf\it\xi}-{\bf\it\psi}}_{2}^{2}$. Using this regularization objective simplifies (5) down to
$\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}=\quantity[\frac{1}{\lambda}\nabla_{{\bf\it\theta}}^{2}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+I_{L}]^{-1},$ (6)
where $I_{L}$ is the $L\crossproduct L$-dimensional identity matrix. Hence the upper-level derivative in (4) can be expressed as
$\begin{split}&\frac{\text{d}F}{\text{d}{\bf\it\theta}_{D}}=\quantity[\frac{1}{\lambda}\nabla_{{\bf\it\theta}}^{2}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+I_{L}]^{-1}\nabla_{{\bf\it\theta}}\mathcal{L}^{D}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{D})+\nabla_{{\bf\it\theta}}\mathcal{L}^{D}\quantity({\bf\it\theta}_{D},{\bf\it\phi}_{D}).\end{split}$ (7)
The inverse Hessian-vector product in (7) is computationally infeasible to calculate directly, so it is approximated using the conjugate gradient (CG) method [47, 48]. We employ a layer-wise implementation of the CG method based on that of [30] and refer to their work for more details on applying CG in a deep learning setup with BLO. While CG has proven to be an effective approach for approximating the inverse Hessian-vector products in previous works [49, 32, 30], it still introduces some computation and memory overhead due to its need for evaluations of multiple Hessian vector products. Nonetheless, the computation time of BiSSL in our experiments is merely a fraction of the self-supervised pretraining time, outlined in Section C.1 of Appendix C. Future work could explore alternative established methods for upper-level gradient approximation [26, 50, 51] that offer greater computational efficiency without compromising downstream task performance.
With an explicit expression of the IJ in (6), we can interpret the impact of the scaling factor $\lambda$ from (3) and (7): When $\lambda$ is very large, the dependence of the lower-level objective on the upper-level parameters ${\bf\it\theta}_{D}$ is also very large. This effectively drives the lower-level backbone parameters toward the trivial solution ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})={\bf\it\theta}_{D}$. Meanwhile, the IJ in (6) approximately equals $I_{L}$, thereby diminishing the influence of the lower-level objective on the upper-level gradient in (7). This roughly makes the task of the upper-level equivalent to conventional fine-tuning. Conversely, if $\lambda$ is very small, the lower-level objective in (3) effectively defaults to conventional pretext task training. Additionally, the implicit jacobian in (6) would consist of numerically tiny entries, making the optimization of the first term in the upper-level objective in (2) equivalent to probing of the downstream head on the frozen pretext backbone ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$.
3.4 Training Algorithm and Pipeline
Algorithm 1 BiSSL Training Algorithm
1:Input: Backbone and pretext head initializations ${\bf\it\theta}$, ${\bf\it\phi}_{P}$, ${\bf\it\phi}_{D}$. Training objectives $\mathcal{L}^{P}$, $\mathcal{L}^{D}$. Optimizers $\text{opt}_{P}$, $\text{opt}_{D}$. Regularization Weight $\lambda\in\mathbb{R}_{+}$. Number of training stage alternations $T\in\mathbb{N}$ with upper and lower-level iterations $N_{U},N_{L}\in\mathbb{N}$.
2:
3:Initialize ${\bf\it\theta}_{P}\leftarrow{\bf\it\theta}$ and ${\bf\it\theta}_{D}\leftarrow{\bf\it\theta}$.
4:
5:for $t=1,\ldots,T$ do
6:     for $n=1,\ldots,N_{L}$ do $\triangleright$ Lower-level
7:         Compute $\mathbf{g}_{{\bf\it\phi}_{P}}=\nabla_{{\bf\it\phi}}\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi})|_{{\bf\it\phi}={\bf\it\phi}_{P}}$
8:         Compute $\mathbf{g}_{{\bf\it\theta}_{P}}=\nabla_{{\bf\it\theta}}\mathcal{L}^{P}\quantity({\bf\it\theta},{\bf\it\phi}_{P})|_{{\bf\it\theta}={\bf\it\theta}_{P}}+\lambda\quantity({\bf\it\theta}_{P}-{\bf\it\theta}_{D})$.
9:         Update ${\bf\it\phi}_{P}\leftarrow\text{opt}_{P}({\bf\it\phi}_{P},\mathbf{g}_{{\bf\it\phi}_{P}})$ and ${\bf\it\theta}_{P}\leftarrow\text{opt}_{P}({\bf\it\theta}_{P},\mathbf{g}_{{\bf\it\theta}_{P}})$.
10:
11:     for $n=1,\ldots,N_{U}$ do $\triangleright$ Upper-level
12:         Compute $\mathbf{g}_{{\bf\it\phi}_{D}}=\nabla_{{\bf\it\phi}}\mathcal{L}^{D}({\bf\it\theta}_{P},{\bf\it\phi})|_{{\bf\it\phi}={\bf\it\phi}_{D}}+\nabla_{{\bf\it\phi}}\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi})|_{{\bf\it\phi}={\bf\it\phi}_{D}}$.
13:         Compute $\mathbf{v}=\nabla_{{\bf\it\theta}}\mathcal{L}^{D}\quantity({\bf\it\theta},{\bf\it\phi}_{D})|_{{\bf\it\theta}={\bf\it\theta}_{P}}$.
14:         Approximate $\mathbf{v}_{\text{IJ}}\approx\quantity[\frac{1}{\lambda}\nabla_{{\bf\it\theta}}^{2}\mathcal{L}^{P}\quantity({\bf\it\theta},{\bf\it\phi}_{P})|_{{\bf\it\theta}={\bf\it\theta}_{P}}+I_{L}]^{-1}\mathbf{v}$. $\triangleright$ Use CG
15:         Compute $\mathbf{g}_{{\bf\it\theta}_{D}}=\mathbf{v}_{\text{IJ}}+\nabla_{{\bf\it\theta}}\mathcal{L}^{D}\quantity({\bf\it\theta},{\bf\it\phi}_{D})|_{{\bf\it\theta}={\bf\it\theta}_{D}}$.
16:         Update ${\bf\it\phi}_{D}\leftarrow\text{opt}_{D}({\bf\it\phi}_{D},\mathbf{g}_{{\bf\it\phi}_{D}})$ and ${\bf\it\theta}_{D}\leftarrow\text{opt}_{D}({\bf\it\theta}_{D},\mathbf{g}_{{\bf\it\theta}_{D}})$.
17:
18:Return: Backbone Parameters $\boldsymbol{\theta}_{P}$.
Algorithm 1 outlines the proposed training algorithm, which iteratively alternates between solving the lower-level (3) and upper-level (2) optimization problems in BiSSL. The lower-level training optimizes the pretext task objective, while additionally including the gradient of the regularization term $r$ for the backbone parameter updates, complying with (3). For the upper-level training, the gradient with respect to the backbone parameters as represented by the left term on the right-hand side in (7), is approximated using the CG method. We remark that allowing $N_{U}>1$ deviates from conventional BLO implementations. However, as documented in Section C.3 of Appendix C, this modification was found to be beneficial for downstream performance. The algorithm is applicable to any common pretext and downstream tasks, with the only supplementary requirement being that the pretext task objective is twice-differentiable.
Figure 1 illustrates the proposed training pipeline with BiSSL alongside the conventional SSL pipeline. Pretext pretraining on the unlabeled dataset $\mathcal{D}^{P}$ provides initializations of ${\bf\it\theta}$ and ${\bf\it\phi}_{P}$, after which the downstream head ${\bf\it\phi}_{D}$ is fitted on top of the frozen backbone ${\bf\it\theta}$ using the downstream dataset $\mathcal{D}^{D}$. BiSSL training is then conducted as outlined in Algorithm 1, yielding an updated backbone ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$, which serves as the initialization for final supervised fine-tuning on the downstream task.
As detailed in Section A.2 in Appendix A, ${\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})$ must satisfy the stationary condition $\nabla_{{\bf\it\theta}}\quantity(\mathcal{L}^{P}({\bf\it\theta},{\bf\it\phi}_{P})+\lambda r({\bf\it\theta}_{D},{\bf\it\theta}))|_{{\bf\it\theta}={\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})}=\mathbf{0}$ to justify the explicit expression of the IJ in (6). This theoretical requirement motivates the placement of BiSSL after the pretraining stage, as a converged initialization ${\bf\it\theta}$ achieved from pretraining ensures the stationary condition is roughly satisfied at the outset. A similar argument applies to ${\bf\it\phi}_{D}$, since a randomly initialized downstream head typically causes rapid updates to ${\bf\it\theta}_{D}$ during early fine-tuning, likely violating the stationary assumption due to its coupling with ${\bf\it\theta}_{P}$ through the regularization objective $r$.
4 Experiments and Results
4.1 Datasets
The ImageNet-1K [36] dataset devoid of labels is used for self-supervised pretraining throughout the main experiments. For downstream fine-tuning and evaluation, we leverage a varied set of natural image classification datasets that encompass a wide array of tasks, including general image classification, fine-grained recognition across species and objects, scene understanding, and texture categorization. The datasets include Food 101 [52], CIFAR10 [53], CIFAR100 [53], Caltech-UCSD Birds-200-2011 (CUB200) [54], SUN397 scene dataset [55], StanfordCars [56], FGVC Aircraft [57], PASCAL VOC 2007 [58], Describable Textures Dataset (DTD) [59], Oxford-IIIT Pets [60], Caltech-101 [61] and Oxford 102 Flowers [62]. All downstream datasets are split into training, validation, and test partitions, with details on how these assignments are made in Section B.1 of Appendix B.
4.2 Implementation Details
4.2.1 Baseline Setup
Pretext Task Training
We conduct experiments using two different types of pretext tasks: SimCLR [1] and Bootstrap Your Own Latent (BYOL) [4]. For SimCLR we use a temperature of $0.5$ and a ResNet-50 [35] backbone. On top of the backbone, a projection head is applied, consisting of two fully connected layers with batch normalization [63] and ReLU [64] followed by a single linear layer. Each layer consists of $256$ neurons. For BYOL, we use constant target decay rate of $0.9995$, with the backbone and projection head architectures identical to those of SimCLR. The additional BYOL-specific prediction head uses an architecture identical to the projection head. The remaining details in this paragraph apply to both SimCLR and BYOL.
[…]
Fine-Tuning on the Downstream Task
[…]
A random grid search of $100$ hyper-parameter configurations for the base learning rates and weight decays is conducted, where one model is fine-tuned for each configuration. Base learning rates and weight decays are log-uniformly sampled over the ranges of $10^{-4}$ to $1$ and $10^{-5}$ to $10^{-2}$, respectively. Validation data accuracy is evaluated after each epoch. The hyper-parameter configuration yielding the best balance between high validation accuracy and low validation loss is considered the optimal hyper-parameter configuration. [^2] In certain scenarios during the experiments, the configuration that achieved the highest validation accuracy also yielded a notably higher relative validation loss. To ensure better generalizability, an alternative configuration with a more favorable trade-off was selected in these cases. The corresponding optimal hyper-parameters for each downstream dataset are documented in Table 3 of Appendix B. For subsequent evaluation on the test data, we train $10$ models with different random seeds, each using the considered optimal hyper-parameter configurations. During the training of each respective model, the model parameters are stored after each epoch if the top-1 validation accuracy (or 11-point mAP for the VOC07 dataset) has increased compared to the previous highest top-1 validation accuracy achieved during training. Top-1 and top-5 test data accuracies (or 11-point mAP for the VOC07 dataset) are evaluated for each of the $10$ models, from which the calculated means and standard deviations of these accuracies are documented.
4.2.2 BiSSL Setup
In this section, we detail each stage of the proposed training pipeline for BiSSL, as outlined in the right part of Figure 1.
Pretext Task
The backbone ${\bf\it\theta}$ and projection head ${\bf\it\phi}_{P}$ are initialized by self-supervised pretraining using a setup identical to the baseline pretext task training setup detailed in Section 4.2.1. This enables reusing the pretrained backbones from the baseline experiments.
Downstream Head Warm-up
The training setup for the downstream head warm-up closely mirrors the fine-tuning setup of Section 4.2.1. The main difference is that only the linear downstream head is fitted on top of the now frozen backbone obtained from the pretext warm-up. Learning rates and weight decays are initially selected based on those listed in Table 3, with adjustments made as needed when preliminary testing indicated a potential for improved convergence. These values are provided in Table 4 in Appendix B. The authors recognize that more optimal hyper-parameter configurations may exist and leave further exploration of this for future refinement. The downstream head warm-up is conducted for $20$ epochs with a constant learning rate.
Lower-level of BiSSL
The training configuration for the lower-level primarily follows the setup described for pretext pretraining in Section 4.2.1, with the modifications outlined here. As specified in (3), the lower-level loss function is the sum of the pretext task objective $\mathcal{L}^{P}$ (e.g. the NT-Xent loss for SimCLR [1]) and the regularization term $r(\boldsymbol{\theta}_{D},\boldsymbol{\theta}_{P})=\frac{1}{2}||{\bf\it\theta}_{D}-{\bf\it\theta}_{P}||_{2}^{2}$, aligning with Algorithm 1. Based on early experiments, the regularization weight $\lambda=0.001$ was selected, as it appeared to strike a well-balanced compromise between the convergence rates of both the lower- and upper-level objectives. The lower-level is trained for the equivalent of approximately $8$ conventional pretraining epochs, with further details provided in the composite configuration paragraph. Accordingly the linear learning rate scheduler warm-up is adjusted to range over $10\cdot N_{L}$ training steps, with $N_{L}$ being the number of lower-level iterations as specified in Algorithm 1.
Upper-level of BiSSL
The upper-level training stage largely mirrors the downstream training setup described in Section 4.2.1, and again, only the differences are addressed here. The weight decays and base learning rates are set to match those obtained from the downstream head warm-up detailed in Table 4 of Appendix B. To approximate the upper-level gradient in  (7), the conjugate gradient method [47, 48] is employed, with implementation details covered in Section B.3 of Appendix B.
Composite Configuration Details of BiSSL
As outlined in Algorithm 1, both lower- and upper-level backbone parameters ${\bf\it\theta}_{P}$ and ${\bf\it\theta}_{D}$ are initialized with the backbone parameters obtained during the pretext warm-up, and the training procedure alternates between solving the lower- and upper-level optimization problems. In this experimental setup, the lower-level performs $N_{L}=20$ gradient steps before alternating to the upper-level, which then conducts $N_{U}=8$ gradient steps. A total of $T=500$ training stage alternations are executed. With the ImageNet dataset and the current batch size of $1024$, there are a total of $1251$ training batches without replacement. Consequently, the $T=500$ training stage alternations correspond to roughly $8$ conventional pretext epochs, a negligible additional training load compared to the $500$ pretext epochs used for the full pretraining process. Additionally, gradient normalization is employed on gradients exceeding $\ell_{2}$-norms of $10$.
Fine-Tuning on the Downstream Task
Subsequent downstream fine-tuning is conducted in a manner identical to that described in the “Fine-Tuning on the Downstream Task” paragraph of section 4.2.1. Table 5 in Appendix B lists the considered optimal hyper-parameter configurations for each dataset.
Table 1: Comparison of classification accuracies between the conventional SSL pipeline and the proposed BiSSL pipeline. Accuracies significantly higher from their counterparts are marked in bold.
Food CIFAR10 CIFAR100 CUB200 SUN397 Cars Aircrafts VOC07 DTD Pets Caltech101 Flowers
SimCLR:
Only FT $75.0\pm 0.1$ $96.1\pm 0.2$ $79.5\pm 0.2$ $49.6\pm 0.3$ $49.3\pm 0.2$ $78.0\pm 0.3$ $52.0\pm 1.0$ $71.0\pm 0.1$ $60.3\pm 0.9$ $73.2\pm 0.3$ $85.7\pm 0.5$ $82.6\pm 0.3$
BiSSL+FT $\mathbf{75.6\pm 0.1}$ $96.4\pm 0.1$ $\mathbf{80.0\pm 0.1}$ $\mathbf{59.2\pm 0.2}$ $\mathbf{51.4\pm 0.2}$ $78.0\pm 0.3$ $\mathbf{55.3\pm 0.5}$ $\mathbf{71.8\pm 0.1}$ $\mathbf{63.4\pm 0.3}$ $\mathbf{79.6\pm 0.3}$ $\mathbf{88.1\pm 0.3}$ $\mathbf{85.6\pm 0.2}$
[…]
BYOL:
Only FT $75.3\pm 0.2$ $96.4\pm 0.1$ $80.1\pm 0.2$ $52.7\pm 0.4$ $47.9\pm 0.3$ $76.9\pm 0.2$ $51.6\pm 0.7$ $69.3\pm 0.1$ $59.5\pm 0.4$ $77.9\pm 0.3$ $86.6\pm 0.4$ $82.2\pm 0.5$
BiSSL+FT $75.4\pm 0.1$ $96.4\pm 0.1$ $\mathbf{80.6\pm 0.1}$ $\mathbf{60.1\pm 0.1}$ $\mathbf{50.2\pm 0.2}$ $\mathbf{78.1\pm 0.3}$ $\mathbf{57.1\pm 0.4}$ $\mathbf{71.1\pm 0.1}$ $\mathbf{62.2\pm 0.1}$ $\mathbf{81.9\pm 0.1}$ $\mathbf{88.3\pm 0.3}$ $\mathbf{86.5\pm 0.1}$
[…]
4.3 Downstream Task Performance
We first assess the impact of using BiSSL compared to the conventional self-supervised training pipeline by evaluating classification accuracies on the various specified downstream datasets. Table 1 report means and standard deviations of top-1 classification accuracies (or the 11-point mAP on the VOC2007 dataset), comparing results obtained from the conventional SSL pipeline with those achieved using the BiSSL pipeline. Table 7 of Section C.2 in Appendix C outlines the corresponding top-5 accuracies. The results reveal BiSSL significantly improves downstream performance on $10$ out of $12$ datasets, with no single result showing a decline in performance compared to the baseline. Moreover, classification accuracies achieved via BiSSL exhibit greater stability, as evidenced by consistently lower or comparable variances relative to the baseline.
Table 2: $\text{AP}_{50}$-scores for object detection on the VOC07+12 dataset using a Faster-RCNN with a R50-C4 backbone.
SimCLR BYOL
Only FT $59.6\pm 0.3$ $54.6\pm 0.5$
BiSSL+FT $\mathbf{60.6\pm 0.1}$ $\mathbf{56.7\pm 0.4}$
Object Detection
To assess if the advantages of BiSSL extend beyond classification tasks, we further assess its performance in the context of object detection. Similarly to the setup of [2, 3], we use a Faster-RCNN with a ResNet-50 C4 backbone [67] on the VOC07+12 [58] dataset and report the standard $\text{AP}_{50}$ metric. We refer to Section B.6 in Appendix B for more implementation details. The results in Table 2 support the previous conclusions, indicating BiSSL significantly improves downstream performance.
Figure 2: Top-1 test classification accuracies on the flowers dataset for separate models pretrained with SimCLR for different durations, comparing the conventional SSL and BiSSL training pipelines. BiSSL consistently achieves significantly higher accuracies than the baseline when the pretraining duration is sufficiently high.
Figure 3: Feature visualizations of BiSSL-trained (upper) vs pretext-only (lower) backbones on the flowers dataset. Colors denote respective classes. Details in Section C.8.
Appendix C provides additional benchmark results. Section C.4 shows that BiSSL also outperforms conventional fine-tuning when using a Masked Autoencoder [7] for self-supervised pretraining. Section C.5 extends comparisons to more baselines, with BiSSL again performing best. Lastly section C.7 presents ImageNet fine-tuning results, where BiSSL performs comparably to conventional fine-tuning, which is expected since pretraining is also conducted on ImageNet, leading to strong alignment between the pretext and downstream distributions.
4.3.1 Varying the pretraining duration
To further assess the robustness of BiSSL, we vary the duration of self-supervised pretraining. Due to computational resource constraints, we adopt a smaller-scale version of the SimCLR setup, using a ResNet-18 backbone and the unlabeled partition of the smaller-scale STL10 [68] dataset for pretraining. Reusing the parameters from Section 4.2.2, BiSSL training corresponds to $100$ conventional pretext epochs using the STL10 dataset. To ensure fair comparison, the pretext-only baselines are accordingly pretrained for $100$ additional epochs. The rest of the setup remains unchanged. We evaluate on the flowers dataset, where BiSSL previously showed notable gains. Figure 3 depicts the final top-1 test accuracies achieved by separate models pretrained for varying durations, with Figure 4 in Appendix C showing the top-5 test accuracies. The results show that BiSSL consistently outperforms the baseline once sufficient pretraining is reached, aligning with the remarks in Section 3.4, suggesting that its advantage stems not from longer training, but rather from the enhanced information sharing it facilitates between the pretext and downstream tasks.
Appendix C presents additional investigative experiments. Section C.1 reports training durations for pretraining and BiSSL, showing that BiSSL requires only a fraction of the computational cost of pretext pretraining. Section C.3 confirms the effectiveness of the BiSSL setup, showing that setting $N_{U}=1$ or excluding the first term of (2) both degrade downstream performance. Section C.6 demonstrates that BiSSL consistently retains is advantage across varying downstream dataset sizes.
4.4 Visual Inspection of Latent Features
To better assess whether BiSSL nudges the latent features toward being more semantically meaningful for downstream tasks, we qualitatively inspect latent spaces using the t-Distributed Stochastic Neighbor Embedding (t-SNE) [69] method. Specifically, we compare features from SimCLR-pretrained backbones to those derived from BiSSL-trained lower-level backbones, as described in Sections 4.2.1 and 4.2.2 respectively. Figure 3 illustrates the results on the flowers dataset, indicating that BiSSL improves downstream feature alignment. Additional plots on a selection of downstream datasets in Section C.8 of Appendix C reinforce this finding, also demonstrating that this trend consistently persists even for datasets where BiSSL did not impose any classification accuracy improvements. Further details regarding the experimental setup are outlined in Section C.8.
5 Conclusion
This study introduces BiSSL, a novel training framework that integrates the self-supervised pretext and downstream fine-tuning objectives into a unified bilevel optimization problem. By explicitly modeling the interdependence between pretraining and fine-tuning, BiSSL facilitates enhanced information sharing between these training stages, yielding model initializations that are better aligned with downstream tasks. We present a task-agnostic training algorithm and pipeline that incorporates BiSSL as an intermediate training stage between the self-supervised pretraining and fine-tuning stages. Experiments across multiple pretext and downstream tasks demonstrate that BiSSL significantly improves downstream performance over the conventional self-supervised learning pipeline. Additional exploratory experiments further support the conclusion that BiSSL enhances downstream task alignment. Overall, BiSSL represents a promising step toward training algorithm designs that better align the self-supervised pretraining and downstream fine-tuning stages via bilevel optimization.
Acknowledgments and Disclosure of Funding
This project is supported by the Pioneer Centre for Artificial Intelligence, Denmark. [^3] https:
www.aicentre.dk
The authors would like to thank Sijia Liu and Yihua Zhang (Michigan State University) providing valuable feedback in a discussion, which helped to refine and solidify our perspective on the topic of integrating bilevel optimization in deep learning.
Appendix A Theoretical Insights and Framework Comparisons in BiSSL
A.1 On the Necessity of the Second Term in the Upper-Level Objective
While the inclusion of the first term $\mathcal{L}^{D}({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\phi}_{D})$ in the upper-level objective in (2) is clearly justified, as it couples the two levels together and enables the upper-level to incorporate information from the lower-level training objective, the second term $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D})$ may at first glance appear redundant. However, this term is crucial for ensuring that the lower-level problem in (3) is non-redundant when using a mean squared difference regularization objective $r({\bf\it\theta}_{D},{\bf\it\theta}_{P})=\frac{1}{2}\norm{{\bf\it\theta}_{D}-{\bf\it\theta}_{P}}_{2}^{2}$, which is the primary regularization objective in our experiments. We here consider the bilevel optimization problem using the aforementioned regularization objective, while also omitting the second term of (2). To simplify notation and focus on the central idea, we omit head parameters in the following derivation, which does not affect the core reasoning.
Let the conventional fine-tuning problem be defined as:
$\min_{{\bf\it\theta}_{D}}\mathcal{L}^{D}({\bf\it\theta}_{D}),$ (8)
and consider the following bilevel optimization problem:
$\underset{\boldsymbol{\theta}_{D}}{\min}\,\,\mathcal{L}^{D}\left(\boldsymbol{\theta}_{P}^{\ast}(\boldsymbol{\theta}_{D})\right)\quad\text{ s.t.}\quad\boldsymbol{\theta}_{P}^{\ast}\left(\boldsymbol{\theta}_{D}\right)\in\underset{\boldsymbol{\theta}_{P}}{\text{argmin}}\,\mathcal{L}^{P}\left(\boldsymbol{\theta}_{P}\right)+\frac{\lambda}{2}\norm{{\bf\it\theta}_{D}-{\bf\it\theta}_{P}}_{2}^{2}.$ (9)
We then have the following result.
Theorem 1: Let $\bar{{\bf\it\theta}}$ be a stationary point of the conventional fine-tuning problem (8). Define:
${\bf\it\theta}_{P}^{\ast}:=\bar{{\bf\it\theta}},\quad{\bf\it\theta}_{D}^{\ast}:=\bar{{\bf\it\theta}}-\frac{1}{\lambda}\nabla_{{\bf\it\theta}}\mathcal{L}^{P}({\bf\it\theta})|_{{\bf\it\theta}=\bar{{\bf\it\theta}}}.$
Then the pair $({\bf\it\theta}_{D}^{\ast},{\bf\it\theta}_{P}^{\ast})$ satisfies the stationary conditions of the bilevel problem (9).
Theorem 1 shows that under the simplified bilevel formulation in Equation (9), any solution $\bar{{\bf\it\theta}}$ to the conventional fine-tuning problem of (8) are recovered by the bilevel optimization in (9). In other words, the bilevel formulation in (9) effectively reduces to the single-level fine-tuning problem.
The flaw of this setup is that the upper-level is solely tasked with influencing the lower-level solution ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$ via ${\bf\it\theta}_{D}$, without any direct incentive to optimize ${\bf\it\theta}_{D}$ itself. Since ${\bf\it\theta}_{D}$ is only directly present in the lower-level, its sole role becomes one of steering the lower-level dynamics such that ${\bf\it\theta}_{P}$ minimizes the upper-level objective, which undermines the point of the bilevel structure.
Including the term $\mathcal{L}^{D}({\bf\it\theta}_{D})$ in the upper-level objective mitigates this issue by enforcing that ${\bf\it\theta}_{D}$ must itself contribute to satisfying the downstream objective, thereby preventing the lower-level solution from being solely steered to fit the upper-level goal. This yields a genuinely bilevel optimization structure, where both ${\bf\it\theta}_{D}$ and ${\bf\it\theta}_{P}$ are subject to interdependent constraints, promoting a more meaningful interplay between the pretext and downstream tasks.
Proof (Theorem 1): By assumption, $\bar{{\bf\it\theta}}$ is a stationary point of the fine-tuning objective in (8), hence it is straightforward to see for the upper-level in (9) that
$\nabla_{{\bf\it\theta}}\mathcal{L}^{D}({\bf\it\theta})|_{{\bf\it\theta}={\bf\it\theta}_{P}^{\ast}}=\nabla_{{\bf\it\theta}}\mathcal{L}^{D}({\bf\it\theta})|_{{\bf\it\theta}=\bar{{\bf\it\theta}}}=\mathbf{0}.$ (10)
Now, defining the lower-level objective as
$G({\bf\it\theta}_{D},{\bf\it\theta}_{P})=\mathcal{L}^{P}\left(\boldsymbol{\theta}_{P}\right)+\frac{\lambda}{2}\norm{{\bf\it\theta}_{D}-{\bf\it\theta}_{P}}_{2}^{2},$ (11)
we similarly evaluate
$\nabla_{{\bf\it\theta}}G({\bf\it\theta}_{D}^{\ast},{\bf\it\theta})|_{{\bf\it\theta}={\bf\it\theta}_{P}^{\ast}}$ $=\nabla_{{\bf\it\theta}}\mathcal{L}^{P}({\bf\it\theta})|_{{\bf\it\theta}={\bf\it\theta}_{P}^{\ast}}+\lambda({\bf\it\theta}_{D}^{\ast}-{\bf\it\theta}_{P}^{\ast})$ (12)
$=\nabla_{{\bf\it\theta}}\mathcal{L}^{P}({\bf\it\theta})|_{\theta=\bar{{\bf\it\theta}}}+\lambda(\bar{{\bf\it\theta}}-\frac{1}{\lambda}\nabla_{{\bf\it\theta}}\mathcal{L}^{P}({\bf\it\theta})|_{\theta=\bar{{\bf\it\theta}}}-\bar{{\bf\it\theta}})$ (13)
$=\mathbf{0}.$ (14)
Hence, the pair $({\bf\it\theta}_{D}^{\ast},{\bf\it\theta}_{P}^{\ast})$ satisfies the stationary conditions of the bilevel problem in (9). ∎
A.2 Derivation of the Implicit Jacobian
Assume the setup of the BiSSL optimization problem described in (2) and (3). In the following derivations, we will assume that ${\bf\it\phi}_{P}$ is fixed, allowing us to simplify the expressions involved. To streamline the notation further, we continue to use the convention $\nabla_{{\bf\it\xi}}h({\bf\it\xi})|_{{\bf\it\xi}={\bf\it\psi}}:=\nabla_{{\bf\it\xi}}h({\bf\it\psi})$, when it is clear from context which variables are differentiated with respect to. Under these circumstances, we then define the lower-level objective from (3) as
$G({\bf\it\theta}_{D},{\bf\it\theta}_{P}):=\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi}_{P})+\lambda r({\bf\it\theta}_{D},{\bf\it\theta}_{P}).$ (15)
Recalling that $r$ is a convex regularization objective, adequate scaling of $\lambda$ effectively “convexifies” the lower-level objective $G$, a strategy also employed on the lower-level objective in previous works [30, 27, 26]. This is advantageous because assuming convexity of $G$ ensures that for any ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$, there exists a corresponding $\hat{{\bf\it\theta}}_{P}\in\mathbb{R}^{L}$ that satisfies the stationary condition $\nabla_{{\bf\it\theta}_{P}}G({\bf\it\theta}_{D},\hat{{\bf\it\theta}}_{P})=\mathbf{0}$. In other words, we are assured that a potential minimizer of $G({\bf\it\theta}_{D},\cdot)$ exists for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$. Now, further assume that the gradient $\nabla_{{\bf\it\theta}_{P}}G({\bf\it\theta}_{D},{\bf\it\theta}_{P})$ is continuously differentiable and that the Hessian matrix $\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},\hat{{\bf\it\theta}}_{P})$ is invertible for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$. Under these conditions, the implicit function theorem [70, 71] guarantees the existence of an implicit unique and differentiable function ${\bf\it\theta}_{P}^{\ast}:\mathcal{N}({\bf\it\theta}_{D})\to\mathbb{R}^{L}$, with $\mathcal{N}({\bf\it\theta}_{D})$ being a neighborhood of ${\bf\it\theta}_{D}$, that satisfies ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})=\hat{{\bf\it\theta}}_{P}$ and $\nabla_{{\bf\it\theta}_{P}}G(\tilde{{\bf\it\theta}}_{D},{\bf\it\theta}_{P}^{\ast}(\tilde{{\bf\it\theta}}_{D}))=\mathbf{0}$ for all $\tilde{{\bf\it\theta}}_{D}\in\mathcal{N}({\bf\it\theta}_{D})$.
As we then conclude that the lower-level solution ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$ is indeed a differentiable function under these conditions, this justifies that the expression
$\frac{\text{d}}{\text{d}{\bf\it\theta}_{D}}\quantity[\nabla_{{\bf\it\theta}_{P}}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))]=\mathbf{0}$
is valid for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$. By applying the chain rule, the expression becomes
$\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))+\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))=\mathbf{0}.$
Recalling that $\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))$ is assumed to be invertible, the implicit jacobian (IJ) $\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}$ can be isolated
$\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}=-\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))\quantity[\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))]^{-1},$
and by substituting the expression for $G$ from (15), the expression becomes
$\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}$ $=-\lambda\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}r({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))\quantity[\nabla_{{\bf\it\theta}_{P}}^{2}\quantity(\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+\lambda r\quantity({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})))]^{-1}$
$=-\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}r({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))\quantity[\nabla_{{\bf\it\theta}_{P}}^{2}\quantity(\frac{1}{\lambda}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+r\quantity({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})))]^{-1}.$ (16)
To summarize, given the following assumptions:
  • The lower-level pretext head parameters ${\bf\it\phi}_{P}$ are fixed.
  • $G$ is convex such that $\nabla_{{\bf\it\theta}_{P}}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))=\mathbf{0}$ is fulfilled for every ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$.
  • The Hessian matrix $\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))$ exists and is invertible for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$.
Then, the IJ $\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}$ can be explicitly expressed by (16). The authors acknowledge that an explicit expression for the IJ without fixing ${\bf\it\phi}_{P}$ is achievable, though this is left for future exploration.
A.3 Distinction from Bilevel Optimization in Meta-Learning
While bilevel optimization (BLO) has been applied in meta-learning frameworks such as MAML [31], Sign-MAML [72] and iMAML [30], BiSSL represents a distinct application and implementation of BLO, tailored for the challenges of self-supervised learning (SSL). In the aforementioned works, BLO is primarily utilized to address few-shot learning scenarios, focusing on efficiently adapting models to new tasks with minimal labeled data. Conversely, BiSSL applies BLO to concurrently manage the more complex task of self-supervised pretext pretraining on unlabeled with downstream fine-tuning on labeled data. Another key distinction is that in meta-learning, the upper- and lower-level objectives are closely related, with the upper-level objective formulated as a summation of the lower-level tasks. In contrast, BiSSL involves fundamentally distinct objectives at each level, utilizing separate datasets and tasks for pretraining and fine-tuning. This design allows BiSSL to better align the pretrained model with the requirements of a specific downstream task. Conversely, the BLO in meta-learning aims to broadly generalize across a wide range of tasks, prioritizing adaptability rather than task-specific optimization. Lastly, unlike BiSSL, the meta-learning frameworks reinitialize the lower-level backbone parameters with a copy of the upper-level parameters at every iteration.
Appendix B Default Experiment Details
B.1 Dataset Partitions
The Caltech-101 [61] dataset does not come with a pre-defined train/test split, so the same convention as previous works is followed [1, 73, 74], where 30 random images per class are selected for the training partition, and the remaining images are assigned for the test partition. For the DTD [59] and SUN397 [55] datasets, which offer multiple proposed train/test partitions, the first splits are used, consistent with the approach in [1].
For downstream hyperparameter optimization, portions of the training partitions from each respective labeled dataset are designated as validation datasets. The FGVC Aircraft [57], Oxford 102 Flowers [62], DTD, and Pascal VOC 2007 [58] datasets already have designated validation partitions. For all the remaining labeled datasets, the validation data partitions are randomly sampled while ensuring that class proportions are maintained. Roughly $20\%$ of the training data is allocated for validation.
B.2 Downstream Task Fine-Tuning of the Baseline Setup
In Table 3, the learning rates and weight decays used for each respective downstream dataset of the experiments described in Section 4.2.1 are outlined.
Table 3: Optimal hyper-parameter configurations used for downstream fine-tuning after conventional pretext pretraining.
Dataset SimCLR BYOL
Learning Rate Weight Decay Learning Rate Weight Decay
Food $0.0167$ $0.00613$ $0.0513$ $0.00147$
CIFAR10 $0.0033$ $0.00158$ $0.0014$ $0.00106$
CIFAR100 $0.0027$ $0.00012$ $0.0023$ $0.00012$
CUB200 $0.0409$ $0.0084$ $0.0095$ $0.00594$
SUN397 $0.0069$ $0.00003$ $0.004$ $0.00003$
Cars $0.0377$ $0.00454$ $0.115$ $0.00257$
Aircrafts $0.0269$ $0.0038$ $0.0119$ $0.00333$
VOC07 $0.0054$ $0.0089$ $0.0032$ $0.00616$
DTD $0.0514$ $0.0011$ $0.0114$ $0.00011$
Pets $0.0378$ $0.00114$ $0.0044$ $0.00779$
Caltech101 $0.0131$ $0.00005$ $0.0069$ $0.00027$
Flowers $0.2178$ $0.00046$ $0.035$ $0.00262$
B.3 Downstream Head Warmup and Upper-level of BiSSL
Table 4 outlines the learning rates and weight decays used for the downstream head warm-up and upper-level of BiSSL of each respective downstream dataset, as described in the BiSSL experimental setup of Section 4.2.2.
Table 4: Hyper-parameters used for the Downstream Head Warm-up and Upper-level of BiSSL.
Dataset SimCLR BYOL
Learning Rate Weight Decay Learning Rate Weight Decay
Food $0.03$ $0.001$ $0.035$ $0.002$
CIFAR10 $0.015$ $0.001$ $0.01$ $0.001$
CIFAR100 $0.01$ $0.0001$ $0.01$ $0.001$
CUB200 $0.03$ $0.001$ $0.015$ $0.0001$
SUN397 $0.015$ $0.00005$ $0.01$ $0.00005$
Cars $0.035$ $0.001$ $0.04$ $0.002$
Aircrafts $0.03$ $0.005$ $0.015$ $0.003$
VOC07 $0.015$ $0.005$ $0.005$ $0.006$
DTD $0.03$ $0.001$ $0.015$ $0.0001$
Pets $0.03$ $0.001$ $0.02$ $0.002$
Caltech101 $0.03$ $0.0001$ $0.015$ $0.0002$
Flowers $0.05$ $0.0005$ $0.035$ $0.002$
The first term of the upper-level gradient (7) is approximated using the Conjugate Gradient (CG) method [47, 48]. Our implementation follows a similar structure to that used in [30], employing $N_{c}=5$ iterations and a dampening term $\lambda_{\text{damp}}=10$. Given matrix $A$ and vector $\mathbf{v}$, the CG method iteratively approximates $A^{-1}\mathbf{v}$, which requires evaluation of multiple matrix-vector products $A\mathbf{d}_{1}$, $\ldots$, $A\mathbf{d}_{N_{c}}$. In practice, storing the matrix $A$ (in our case, the Hessian $\nabla^{2}_{{\bf\it\theta}_{P}}\mathcal{L}^{P}({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})$) in its full form is often infeasible. Instead, a function that efficiently computes the required matrix-vector products instead of explicitly storing the matrix is typically utilized. For transparency, the function employed in our setup is detailed in Algorithm 2. This approach ensures that the output of the CG algorithm is an approximation of the inverse Hessian-vector product in the first term of Equation (7) as intended.
Algorithm 2 Hessian Vector Product Calculation $f_{H}$ (To use in the CG Algorithm)
1:Input: Input vector $\mathbf{v}$. Model parameters ${\bf\it\theta}_{P}$, ${\bf\it\phi}_{P}$. Training objective $\mathcal{L}^{P}$. Lower-level data batche $\mathbf{x}$. Regularization weight $\lambda$ and dampening $\lambda_{\text{damp}}$.
2:
3:$\pi({\bf\it\theta}_{P})\leftarrow\quantity(\nabla_{\boldsymbol{\theta}}\mathcal{L}^{P}\left(\boldsymbol{\theta},\boldsymbol{\phi}_{P};\mathbf{x}\right)|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{P}})^{T}\mathbf{v}$
4:$\mathbf{g}\leftarrow\nabla_{{\bf\it\theta}}\pi({\bf\it\theta})\big{|}_{\boldsymbol{\theta}=\boldsymbol{\theta}_{P}}$$\triangleright$ Memory efficient calculation of $\nabla_{{\bf\it\theta}}^{2}\mathcal{L}^{P}({\bf\it\theta},{\bf\it\phi}_{P};\mathbf{x})|_{{\bf\it\theta}={\bf\it\theta}_{P}}\mathbf{v}$.
5:$\mathbf{y}\leftarrow\mathbf{v}+\frac{1}{\lambda+\lambda_{\text{damp}}}\mathbf{y}$
6:
7:Return: $f_{H}(\mathbf{v}):=\mathbf{y}$
B.4 Composite Configuration of BiSSL
To avoid data being reshuffled between every training stage alternation, the respective batched lower- and upper-level training datasets are stored in separate stacks from which data is drawn. The stacks are only “reset” when the number of remaining batches is smaller than the number of gradient steps required before alternating to the other level. For example, the lower-level stack is reshuffled every fourth training stage alternation. If the downstream dataset does not provide enough data for making $N_{U}=8$ batches with non-overlapping data points, the data is simply reshuffled every time the remaining number of data points is smaller than the upper-level batch size (e.g. $256$ images in the classification experiments).
B.5 Downstream Fine-Tuning after BiSSL
The learning rates and weight decays used for downstream fine-tuning after BiSSL for each respective downstream dataset are outlined in Table 5. Section 4.2.2 outlines the experimental setup.
Table 5: Optimal hyper-parameter configurations used for downstream fine-tuning after BiSSL.
Dataset SimCLR BYOL
Learning Rate Weight Decay Learning Rate Weight Decay
Food $0.0037$ $0.00018$ $0.0496$ $0.00096$
CIFAR10 $0.001$ $0.00278$ $0.0018$ $0.00014$
CIFAR100 $0.0019$ $0.00002$ $0.0008$ $0.0013$
CUB200 $0.0005$ $0.00331$ $0.0007$ $0.00947$
SUN397 $0.0015$ $0.00001$ $0.0009$ $0.00062$
Cars $0.0289$ $0.0075$ $0.0316$ $0.00632$
Aircrafts $0.0002$ $0.0059$ $0.0003$ $0.00001$
VOC07 $0.0002$ $0.00019$ $0.0004$ $0.0003$
DTD $0.0003$ $0.00002$ $0.0013$ $0.0021$
Pets $0.0014$ $0.00005$ $0.0002$ $0.00039$
Caltech101 $0.0016$ $0.00042$ $0.0005$ $0.00022$
Flowers $0.0189$ $0.00143$ $0.0057$ $0.00207$
B.6 Object Detection
The experimental setup primarily follows the main implementation specified in Section 4.2, with minor modifications specified here.
We utilize a Faster R-CNN with a ResNet-50 C4 backbone [67] as the downstream model. The downstream data augmentations involve rescaling the images so that their longest edge is between196196196196 and $320$ pixels, followed by normalization. Downstream fine-tuning is conducted for $50$ epochs with a batch size of $16$. The random hyperparameter grid search is performed across $50$ distinct configurations of learning rates and weight decays, sampled log-uniformly within the intervals $10^{-4}$ to $10^{-1}$ and $10^{-6}$ to $10^{-2}$, respectively.
Important for the context of BiSSL, we extend the pretext head parameters ${\bf\it\phi}_{P}$ to further include the latter layers of the ResNet-50 that are not part of the ResNet-50 C4 backbone. Lastly, the linear head warm-up is conducted for 5 epochs.
Appendix C Additional Results
C.1 Computation Times
We report the training times required for both pretext pretraining and BiSSL under the experimental setup described in Section 4.2.1 and 4.2.2. SimCLR pretraining was conducted on 8x A100 GPUs, while BYOL used 4x A100 GPUs. BiSSL training was measured on the DTD dataset using 4x A40 GPUs. Since the number of gradient steps is kept constant across all datasets in the main experiments, this measurement is representative of BiSSL’s computational cost across other datasets as well.
The reported BiSSL training time reflects the total training time required on a single GPU, therefore we adjust for multi-GPU usage (e.g., for BiSSL, we multiply the total time taken by 4, as we used 4 GPUs). The results are documented in Table 6, and we observe that BiSSL requires only a small fraction of the computational cost of pretraining.
Table 6: Total GPU hours required for pretext pretraining and BiSSL training. BiSSL requires only fraction of the computational cost of pretext pretraining.
SimCLR BYOL
Pretext Training $1464$ GPU Hours $1380$ GPU Hours
BiSSL Training $68$ GPU Hours $70$ GPU Hours
C.2 Top-5 Classification Accuracies
Table 7 outlines the corresponding top-5 accuracies of the experiments described in Section 4.3, which re-emphasizes the performance improvements imposed by BiSSL as initially implied by Table 1. Likewise, Figure 4 outlines the corresponding top-5 accuracies of the experiments described in Section 4.3.1.
Table 7: Comparison of top-5 classification accuracies between the conventional SSL pipeline and our proposed BiSSL pipeline. Accuracies that are significantly higher from their counterparts are marked in bold font. Table 1 outlines the top-1 accuracies (and 11-point mAP for the VOC07 dataset).
Food CIFAR10 CIFAR100 CUB200 SUN397 Cars Aircrafts DTD Pets Caltech101 Flowers
SimCLR:
Only FT $91.3\pm 0.1$ $100.0\pm 0.0$ $96.4\pm 0.1$ $75.2\pm 0.4$ $79.4\pm 0.4$ $93.7\pm 0.2$ $81.8\pm 0.5$ $85.8\pm 0.6$ $94.8\pm 0.4$ $98.1\pm 0.1$ $94.5\pm 0.2$
BiSSL+FT $\mathbf{93.1\pm 0.1}$ $100.0\pm 0.0$ $\mathbf{96.9\pm 0.1}$ $\mathbf{86.5\pm 0.2}$ $\mathbf{81.5\pm 0.1}$ $93.7\pm 0.1$ $\mathbf{85.0\pm 0.2}$ $\mathbf{88.7\pm 0.3}$ $\mathbf{96.5\pm 0.1}$ $\mathbf{98.7\pm 0.1}$ $\mathbf{95.8\pm 0.1}$
[…]
BYOL:
Only FT $91.8\pm 0.1$ $100.0\pm 0.0$ $96.3\pm 0.1$ $77.5\pm 0.4$ $77.8\pm 0.3$ $93.9\pm 0.2$ $81.8\pm 0.6$ $84.9\pm 0.4$ $95.6\pm 0.2$ $98.0\pm 0.1$ $93.6\pm 0.3$
BiSSL+FT $91.8\pm 0.2$ $100.0\pm 0.0$ $\mathbf{96.7\pm 0.1}$ $\mathbf{86.7\pm 0.2}$ $\mathbf{80.1\pm 0.2}$ $94.1\pm 0.2$ $\mathbf{87.5\pm 0.3}$ $\mathbf{87.7\pm 0.3}$ $\mathbf{97.2\pm 0.1}$ $\mathbf{98.7\pm 0.1}$ $\mathbf{95.7\pm 0.1}$
[…]
Figure 4: Top-5 test classification accuracies on the Flowers dataset for separate models pretrained with SimCLR for different durations, comparing the conventional and BiSSL training pipelines. The corresponding top-1 accuracies are shown in Figure 3.
C.3 Impact of Upper-Level Iterations and the IJ Term
This section presents ablation studies showcasing how certain modifications in the BiSSL upper-level optimization affect downstream classification performance. We specifically modify the number of upper-level iterations and the inclusion of the implicit Jacobian (IJ) term in (2).
Due to resource constraints, we evaluate these variations using a representative subset of the 12 datasets from the main experiments (Section 4.1). Specifically, we chose the Pets [60], DTD [59], VOC07 [58], and Flowers [62] datasets, which together cover a diverse range of image domains. Pets and Flowers contain fine-grained object categories, VOC07 features multi-label natural scenes, and DTD consists of texture-based inputs.
All experiments in this section use SimCLR as the pretext task.
Upper-Level Iterations
BiSSL deviates from conventional bilevel optimization implementations by allowing multiple upper-level gradient updates before alternating back to the lower level, controlled via the hyperparameter $N_{U}$ in Algorithm 1. We claim that conducting additional upper-level iterations improves convergence efficiency during training.
To evaluate this design choice, we conduct experiments with $N_{U}=1$ to match conventional bilevel update schedules, while keeping all other aspects identical to the default BiSSL setup in Section 4.2.2, aside from performing the subsequent fine-tuning hyperparameter grid search on 50 different hyperparameter combinations. Since $N_{U}=1$ results in fewer total upper-level updates, we also experiment with a longer training duration by increasing the number of training stage alternations from $T=500$ to $T=4000$, matching the total number of gradient updates conducted in the default setup. Notice however that this results in significantly more number of total lower-level steps ($20\cdot 4000=80,000$) conducted compared to the default setting ($40\cdot 500=20,000$). Despite this, the results in the two middle rows of Table 8 show that neither of the $N_{U}=1$ variants outperforms the default configuration in the top row with $N_{U}=8$ and $T=500$.
Discarding the IJ Term
We further examine the importance of the first term in the upper-level objective of Equation (2) that includes the IJ, which accounts for the influence of lower-level updates on the downstream loss. To do so, we discard this term and retain only the downstream loss $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D})$ in the upper-level objective. The rest of the setup follows the default configuration in Section 4.2.2, except that the fine-tuning hyperparameter grid search was conducted for 50 different combinations. As shown in the final row of Table 8, removing the IJ term leads to notable drops in downstream accuracy, underscoring its role in guiding the BiSSL optimization toward better downstream task alignment.
Table 8: Ablation results showing the impact on downstream classification accuracy from setting $N_{U}=1$ and from removing the first term of the upper-level objective in Equation (2). All experiments use SimCLR as the pretext task. The top row ($N_{U}=8,T=500$) corresponds to the default BiSSL configuration described in Section 4.2.2. The four datasets were selected based on input diversity and computational resource constraints.
BiSSL+FT Pets DTD VOC07 Flowers
$N_{U}=8,T=500$ $79.6\pm 0.3$ $63.4\pm 0.4$ $71.8\pm 0.1$ $85.6\pm 0.2$
$N_{U}=1,T=500$ $75.2\pm 0.4$ $61.5\pm 0.5$ $71.2\pm 0.1$ $82.2\pm 0.3$
$N_{U}=1,T=4000$ $77.6\pm 0.4$ $63.1\pm 0.3$ $71.2\pm 0.1$ $84.4\pm 0.2$
IJ Discarded $75.8\pm 0.4$ $62.9\pm 0.2$ $71.8\pm 0.1$ $82.0\pm 0.3$
C.4 Masked Autoencoder
Our main experiments demonstrate the effectiveness of BiSSL using contrastive pretext tasks and ResNet-based architectures. We here evaluate BiSSL in a slightly different setting using a Masked Autoencoder (MAE) [7].
Table 9: Downstream classification performance when using a MAE as the pretext task. BiSSL improves performance and reduces variance across all four datasets. The four datasets were selected based on input diversity and computational resource constraints.
MAE Pets DTD VOC07 Flowers
Only FT $81.3\pm 0.5$ $61.3\pm 0.7$ $69.5\pm 0.3$ $81.1\pm 1.1$
BiSSL+FT $\mathbf{83.2\pm 0.2}$ $\mathbf{63.5\pm 0.4}$ $\mathbf{71.8\pm 0.2}$ $\mathbf{87.2\pm 0.3}$
[…]
We evaluate performance on a representative subset of the twelve downstream datasets used in the main experiments, due to resource limitations: Pets [60], DTD [59], VOC07 [58], and Flowers [62]. These datasets were selected to represent a broad range of data distributions and classification challenges, from fine-grained object categories (Pets, Flowers), to texture-based recognition (DTD), and multi-label natural scenes (VOC07).
The results in Table 9 show that BiSSL again yields consistent improvements across all datasets and reduces variance in downstream accuracy compared to solely conducting conventional fine-tuning.
C.5 Additional Baseline Comparisons
This section introduces two additional baseline methods that involve intermediate modifications to the pretrained backbone prior to fine-tuning, aiming to provide further context for the effectiveness of BiSSL.
Due to computational constraints, we evaluate these baselines on a subset of four datasets from the main experiments, namely the Pets [60], DTD [59], VOC07 [58], and Flowers [62] datasets. These datasets were selected to represent distinct challenges and input domains in visual recognition, including fine-grained classification (Pets, Flowers), texture-based classification (DTD), and multi-label object recognition in natural scenes (VOC07).
In all experiments within this section, we use SimCLR as the pretext task.
Table 10: Top-1 and Top-5 accuracies on the DTD dataset using different values of $w$ for the weighted objective baseline in (17), compared with conventional fine-tuning and BiSSL. Performance peaks at $w=0.25$, but still lags behind BiSSL.
Accuracy $w=\mathbf{0}$ $w=\mathbf{0.05}$ $w=\mathbf{0.1}$ $w=\mathbf{0.25}$ $w=\mathbf{0.5}$ $w=\mathbf{0.75}$ Only FT BiSSL+FT
Top-1 $60.1\pm 0.3$ $61.6\pm 0.2$ $61.7\pm 0.2$ $\underline{62.2\pm 0.3}$ $60.8\pm 0.3$ $60.3\pm 0.5$ $60.3\pm 0.9$ $\mathbf{63.4\pm 0.3}$
Top-5 $\underline{86.3\pm 0.3}$ $\underline{86.2\pm 0.3}$ $\underline{86.2\pm 0.3}$ $\underline{86.1\pm 0.3}$ $\underline{86.2\pm 0.3}$ $85.7\pm 0.3$ $85.8\pm 0.6$ $\mathbf{88.7\pm 0.3}$
Table 11: Comparison of downstream classification accuracy using alternative baseline methods: a weighted combination of pretext and downstream objectives, and NoisyTune. All experiments use SimCLR as the pretext task. The four datasets were selected based on input diversity and computational resource constraints. BiSSL significantly outperforms all baselines.
Pets DTD VOC07 Flowers
Only FT $73.2\pm 0.3$ $60.3\pm 0.9$ $71.0\pm 0.1$ $82.6\pm 0.3$
PretextDownstreamSum+FT $76.4\pm 0.4$ $62.2\pm 0.3$ $67.1\pm 0.1$ $83.8\pm 0.1$
NoisyTune+FT $73.4\pm 0.3$ $60.0\pm 0.7$ $70.3\pm 0.2$ $82.6\pm 0.2$
BiSSL+FT  $\mathbf{79.6\pm 0.3}$ $\mathbf{63.4\pm 0.4}$ $\mathbf{71.8\pm 0.1}$ $\mathbf{85.6\pm 0.2}$
Weighted Sum of Pretext and Downstream Objectives
The first baseline attempts a simple alignment of the pretext and downstream tasks by solving the single-level optimization problem:
$\min_{{\bf\it\theta},{\bf\it\phi}_{P},{\bf\it\phi}_{D}}(1-w)\mathcal{L}^{P}({\bf\it\theta},{\bf\it\phi}_{P};\mathcal{D}^{P})+w\mathcal{L}^{D}({\bf\it\theta},{\bf\it\phi}_{D};\mathcal{D}^{D}),$ (17)
where $w\in[0,1]$ controls the relative weighting of the pretext and downstream objectives. While this method offers a more straightforward alternative to BiSSL, it lacks the principled bilevel optimization structure, and as our results will show, its performance improvements are less effective.
We use the lower-level optimizer configuration from BiSSL (see Section 4.2.2) and train for $4000$ steps, matching the total number of upper-level iterations in the default BiSSL configuration. Fine-tuning follows the main setup but with a grid search over 50 hyperparameter combinations. Backbone gradients for the two loss terms are computed separately on their respective distinct mini-batches, scaled by $1-w$ and $w$ respectively, and then summed prior to the update step.
To determine the suitable size of $w$, we conducted a sweep on the DTD dataset over different values of $w$ in the range between 00 and $1$, evaluating both top-1 and top-5 classification performance. For each value of $w$, we conduct the subsequent fine-tuning hyperparameter grid search over 25 different combinations. The results of the sweep are shown in Table 10.
The results imply that $w=0.25$ provides the best top-1 classification accuracy. In contrast to BiSSL, the top-5 accuracy does generally not show a significant increase compared to conventional fine-tuning, regardless of the value of $w$. Setting $w=0$ essentially reduces the training to pretext-only learning, reflected by comparable performance to conventional fine-tuning (the "Only FT" column), with slightly lower top-1 accuracy and slightly higher top-5 accuracy. Based on the results, we consider $w=0.25$ as the optimal value for this baseline.
Using $w=0.25$, we summarize the baseline experiments in the "PretextDownstreamSum" row in Table 11, where BiSSL significantly outperforms this baseline across all four datasets.
NoisyTune
We also evaluate NoisyTune [44], which applies Gaussian noise to the backbone weights prior to fine-tuning. We use a noise scale of $0.015$, and conduct a hyperparameter grid search over $50$ different combinations. Results are included in the "NoisyTune" row of Table 11, where BiSSL again demonstrates superior accuracy across all datasets.
C.6 Varying the Downstream Dataset Size
Our downstream datasets already vary widely in size, ranging from $1,020$ to $60,600$ samples, with Table 1 showing no clear correlation between dataset size and accuracy. To further investigate whether this trend persists when considering a single dataset, we conducted additional experiments on the Pets [60] dataset, varying the number of samples per class. The SimCLR pretext task [1] was used, and the fine-tuning hyperparameter grid search was conducted over $25$ combinations, while all other settings were maintained as in the original setup described in Section 4.2.2. We evaluated accuracy using $5$, $10$, $25$, $50$, $75$, samples per class.
Table 12: Varying the Number of Samples per Class on the Pets Dataset. SimCLR is used as the pretext task.
Samples Per Class $\mathbf{5}$ $\mathbf{10}$ $\mathbf{25}$ $\mathbf{50}$ $\mathbf{75}$
Only FT $33.4\pm 0.5$ $45.4\pm 0.5$ $62.5\pm 0.6$ $70.5\pm 0.5$ $73.1\pm 0.2$
BiSSL+FT $\mathbf{37.2\pm 0.6}$ $\mathbf{52.3\pm 0.3}$ $\mathbf{67.2\pm 0.2}$ $\mathbf{73.0\pm 0.3}$ $\mathbf{77.7\pm 0.1}$
[…]
Results are presented in Table 12, which show that BiSSL consistently outperforms conventional fine-tuning, regardless of the number of samples available per class.
C.7 Fine-Tuning on ImageNet
We expect the gains from BiSSL on ImageNet classification [36] as a downstream task to be limited in this specific setting, as pretraining is also conducted on ImageNet, which minimizes the distribution discrepancy between the pretext and downstream tasks. Nonetheless, for the sake of transparency, we conducted fine-tuning experiments on ImageNet within the bounds of our available computational resources.
Our setup consists of $50$ fine-tuning epochs, using a hyperparameter grid search over $25$ configurations. For BiSSL, we use $T=100$ training stage alternations, while keeping the rest of the experimental setup consistent with the description in Section 4.2.2. We use SimCLR as the pretext task. Following standard practice in SSL literature for ImageNet fine-tuning [1, 4, 2], we report the top-1 validation accuracy on models trained with the $1\%$ and $10\%$ subsets using the official splits from [1]. Additionally, we also include results on the full ImageNet dataset ($100\%$). Table 13 shows the results of these experiments.
Table 13: Top-1 validation classification accuracy on ImageNet using $1\%$, $10\%$, and $100\%$ of the labeled training data. SimCLR is used as the pretext task.
ImageNet 1% ImageNet 10% ImageNet 100%
Only FT $29.7\pm 0.1$ $50.9\pm 0.1$ $68.2\pm 0.1$
BiSSL+FT $\mathbf{31.8\pm 0.1}$ $51.1\pm 0.1$ $68.2\pm 0.1$
[…]
As expected, BiSSL yields only modest improvements on ImageNet. The most notable gain appears in the $1\%$ setting, which may be attributed to a greater distribution mismatch between the small labeled subset and the pretraining dataset. This represents the type of scenario where BiSSL is particularly effective at improving alignment. In contrast, the $10\%$ subset is more representative of the full dataset, and the relative advantage of BiSSL correspondingly diminishes.
C.8 Visual inspection of Latent Features
Test data features of the downstream test data processed by backbones trained through conventional pretext pretraining with SimCLR are compared against those additionally trained with BiSSL. This allows for an inspection of the learned representations prior to the final fine-tuning stage.
During the evaluation, it is important to note that the batch normalization layers [63] of the pretrained backbones utilize the running mean and variance inferred during training. Since these pretrained backbones have not been exposed to the downstream datasets during training, their batch normalization statistics may not be optimal for these new datasets. To address this, the downstream training dataset is divided into batches of $256$ samples, and roughly $100$ batches are then forward-passed through the backbones. This procedure ensures that the batch normalization statistics are better suited to the downstream datasets, thereby providing a fairer comparison of the learned representations.
For the dimensionality reduction and visualization of these latent features, the t-Distributed Stochastic Neighbor Embedding (t-SNE) [69] technique is employed. This method allows us to visually assess the clustering and separation of features in the latent space, providing qualitative insights into the semantic structure of the representations learned through BiSSL.
Figures 3 and 5 to 10 illustrate the outcomes of these visual inspections on a selection of the downstream datasets described in Section 4.1, highlighting the differences in feature representations between conventional pretext pretraining and BiSSL.
Figure 5: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the CIFAR10 [53] dataset.
Figure 6: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the CUB200 [54] dataset.
Figure 7: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the Caltech-101 [61] dataset.
Figure 8: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the Pets [60] dataset.
Figure 9: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the Aircrafts [57] dataset.
Figure 10: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the Cars [56] dataset.
Appendix D Discussion of Limitations
In the interest of transparency, we highlight limitations of our current approach that may inform future work.
Our bilevel optimization procedure employs a conjugate gradient (CG) solver, selected for its relative simplicity and prior success in related settings. While this solver proved effective in our experiments, it may not be the most optimal in terms of computational efficiency or compatibility with the non-convex landscape inherent in deep neural networks. Future work could explore alternative solvers or reformulations of the optimization problem to further improve both accuracy and efficiency.
Access to the pretext task, pretraining data (or potentially a subset thereof), and associated pretext head parameters are current requirements for applying BiSSL. This may limit applicability in some constrained environments, although future work could investigate strategies to partially decouple BiSSL from these dependencies.
BiSSL also requires the backbone architectures used in the pretext and downstream stages to be identical. While this is standard in many transfer learning pipelines, it may limit flexibility in scenarios involving architectural modifications, such as the use of parameter-efficient fine-tuning methods [78]. However, this constraint could be mitigated by applying the same architectural modifications consistently across both stages of the BiSSL framework, another avenue worth exploring in future work.
Finally, the benefits from BiSSL are expected to be most pronounced when there is a meaningful distributional discrepancy between the pretraining and downstream tasks. As demonstrated in Section C.7, when the downstream data distribution closely aligns with the pretext task, BiSSL may offer limited improvement. In practice, some degree of distributional shift is often present, but users should be mindful that BiSSL’s effectiveness may depend on the extent of that mismatch.